# Armijo Line-search Makes (Stochastic) Gradient Descent Go Fast

### v1

@@ 5 GD-LS for Non-convex Losses
In this section, we consider non-convex losses that satisfy two alternative gradient domination conditions which enable convergence to the global optimum. In Section 5.1, we analyze the convergence of GD-LS for objectives corresponding to the softmax policy gradient in reinforcement learning. In Section 5.2, we consider the standard Polyak-Łojasiewicz (PL) condition (Karimi et al., 2016; Polyak, 1987) for generalized linear models with a logistic link.
@@
$f$ satisfies a non-uniform gradient domination condition with constant $\zeta\geq 1$, if there exists a $\mu(\theta)>0$ s.t. for all $\theta$,
@@
Gradient domination or Łojasiewicz conditions are satisfied for matrix factorization (Ward & Kolda, 2023), policy gradient in reinforcement learning (Mei et al., 2020) and generalized linear models (Mei et al., 2021). This property has been exploited to prove global convergence guarantees for first-order methods (Karimi et al., 2016; Mei et al., 2021).
@@ 5.1 Gradient domination with $\zeta=1$
We use softmax policy optimization for multi-armed bandits (MAB) as a concrete example that satisfies Assumptions 1, 2 and 3 and Assumption 4 with $\zeta=1$. In particular, we consider an MAB problem with deterministic, known rewards that is often used as a testbed to evaluate policy gradient methods (Xiao, 2022; Mei et al., 2020; Lu et al., 2024). We prove the following proposition in Appendix A.
@@
where $a^{*}:=\operatorname*{arg\,max}_{a\in[K]}r(a)$ is the optimal arm. $f(\theta)$ is non-negative, satisfies Assumption 2 with $L_{0}=0$ and $L_{1}=\frac{3\sqrt{2}}{\Delta}$ssumption 3 with $\nu=\frac{\sqrt{2}}{\Delta}$ and $\omega=0$ and Assumption 4 with $\zeta=1$ and $\mu(\theta)=\pi_{\theta}(a^{*})$. Here, $\Delta:=\max_{a\neq a^{*}}r(a^{*})-r(a)$ is the reward gap and quantifies the problem difficulty.
Softmax policy gradient methods (Williams, 1992) optimize the above non-convex objective using gradient descent, and have been analyzed recently (Mei et al., 2020; Agarwal et al., 2021). We aim to use GD-LS to optimize the objective defined in Proposition 3 for which the GD update is given by: $\theta_{t+1}=\theta_{t}-{\eta_{t}}\,\nabla f(\theta_{t})=\theta_{t}+{\eta_{t}}\nabla_{\theta}\langle\pi_{\theta},r\rangle$, and the Armijo condition is equal to:
$\langle\pi_{\theta_{t+1}},r\rangle$ $\geq\langle\pi_{\theta},r\rangle+c{\eta_{t}}\,\left\|\nabla_{\theta}\langle\pi_{\theta},r\rangle\right\|_{2}^{2}$
Hence, unlike in Lu et al. (2024), implementing GD-LS does not require knowledge of $r(a^{*})$, and hence the identity of the optimal arm. In Proposition 6 in Appendix A, we show that, under additional assumptions, the softmax policy gradient objective for tabular Markov decision processes (MDPs) also satisfies the Assumptions 1, 2, 3 and 4 with $L_{0}=0$ and $\omega=0$. Consequently, we characterize the convergence rate of GD-LS on such problems.
@@
For a fixed $\epsilon>0$, assuming $f(\theta)$ satisfies Assumptions 1, 2 and 3 with $L_{0}=0,\,\omega=0$ and Assumption 4 with $\zeta=1$, GD-LS with $\eta_{\max}=\infty$, requires
@@
iterations to ensure $f(\theta_{T})\leq\epsilon$ where $\mu:=\min_{t\in[T]}\mu(\theta_{t})$.
In order to better understand the implications of Corollary 3, we instantiate the above result for MAB and obtain the following corollary.
@@
For an MAB problem with $K$ arms, rewards bounded in $[0,1]$ and reward gap equal to $\Delta$, GD-LS with a uniform initialization i.e. $\pi_{\theta_{0}}(a)=\frac{1}{K}$ for all $a$, $c=\frac{1}{2}$, $\eta_{\max}=\infty$ requires $T\geq\frac{12^{2}\,K^{2}}{\Delta}\,\ln\left(\frac{1}{\epsilon}\right)$ iterations to guarantee $\langle\pi_{\theta_{T}},r\rangle\geq r(a^{*})-\epsilon$.
Hence, for MAB problems, GD-LS converges at a linear rate. Under additional assumptions, a similar result also holds for tabular MDPs. This is in contrast to GD(1/L) which can only attain an $\Omega\left(\frac{1}{\epsilon}\right)$ convergence rate for both bandits and MDPs (Mei et al., 2020, Theorem 9, 10). The convergence rate of GD-LS matches that of algorithms explicitly designed for this problem, including GD with a specific line-search that requires knowledge of $r(a^{*})$ (Lu et al., 2024), GD with specific increasing step-sizes (Liu et al., 2024), normalized GD (Mei et al., 2021), natural policy gradient (Xiao, 2022) and mirror descent with a log-sum-exp mirror map (Asad et al., 2024).
We note that Assumptions 1, 2 and 3 and Assumption 4 with $\zeta=1$ and $f^{*}=0$ are also satisfied when using the (i) exponential loss to train (ii) two-layer neural networks with a smoothed leaky-ReLU non-linearity and (iii) assuming that the training data is linearly separable (Taheri & Thrampoulidis, 2023, Lemmas 3,5). In this setting, Theorem 1 in Taheri & Thrampoulidis (2023) shows that normalized GD can result in linear convergence for the resulting non-convex objective. Since this problem also satisfies the required assumptions for Corollary 3, we conclude that GD-LS also converges linearly and unlike normalized GD, it does not require knowing problem-dependent constants. Hence, our results demonstrate the universality of GD-LS.
@@ 5.2 Gradient domination with $\zeta=2$
We use generalized linear models (GLM) with a logistic link function as an example. In Proposition 2, we have seen that the objective satisfies Assumptions 1, 2 and 3 with non-zero $L_{0},L_{1}$, $\nu,\omega$. It also satisfies Assumption 4 with $\zeta=2$.
@@
If $\sigma(\cdot)$ is the sigmoid function and $\pi_{i}(\theta):=\sigma(\langle x_{i},\theta\rangle)$, assuming that for all $i\in[n]$, $\left\|x\right\|_{i}\leq 1$, $y_{i}=\pi_{i}(\theta^{*})$ such that $\left\|\theta^{*}\right\|\leq D<\infty$ and $\upsilon(\theta):=\min_{i\in[n]}{\left\{\pi_{i}(\theta)\cdot\left(1-\pi_{i}(\theta)\right)\right\}}$, then the GLM objective in Eq. 4 satisfies Assumption 4 with $\zeta=2$ and $\mu(\theta)=64\,[\upsilon(\theta)]^{2}\,[\min\{\upsilon(\theta),\upsilon(\theta^{*})\}]^{2}$.
Similar to logistic regression, the PL constant $\mu$ depends on $\upsilon(\theta)$. However, unlike logistic regression where $y_{i}\in\{0,1\}$ and $\left\|\theta^{*}\right\|$ can be infinite on separable data, for GLMs, $y_{i}\in(0,1)$ and $\left\|\theta^{*}\right\|$ is bounded. Consequently, $\mu(\theta^{*})>0$ and as long as $\left\|\theta_{t}\right\|<\infty$ for any iterate $t\in[T]$, $\mu(\theta)$ is bounded away from zero. However, we note that this does not preclude the case where the iterates initially diverge away from the solution, resulting in large $\left\|\theta_{t}\right\|$ and small (but non-zero) $\mu(\theta)$. Furthermore, the realizability condition in Lemma 2 implies that $f^{*}=0$. Given these considerations, we prove the following theorem in Appendix D.
@@
For a fixed $\epsilon\in\left(0,\frac{\lambda_{0}}{\lambda_{1}}\right)$, if $f$ satisfies Assumptions 1, 2 and 3 and Assumption 4 with $\zeta=2$ with $f^{*}=0$ and if $\mu:=\min_{t\in[T]}\mu(\theta_{t})$, GD-LS with $\eta_{\max}=\infty$, requires
@@
Proof Sketch.
Using the gradient domination condition with the Armijo condition in Eq. 6 and the lower-bound on ${\eta_{t}}$ from Lemma 1, we get that, $f(\theta_{t+1})\leq f(\theta_{t})-\frac{\mu\,f(\theta_{t})}{\lambda_{0}+\lambda_{1}\,f(\theta_{t})}$. For the first $\tau$ iterations, we have that $f(\theta_{t})\geq\frac{\lambda_{0}}{\lambda_{1}}$. Since $\lambda_{0}\leq\lambda_{1}\,f(\theta_{t})$, by recursing for $\tau$ iterations, we conclude that $f(\theta_{\tau})\leq f(\theta_{0})-\tau\,\frac{\mu}{2\,\lambda_{1}}$, meaning that $\tau\leq\frac{2\lambda_{1}\,f(\theta_{0})}{\mu}$. This corresponds to Phase 1. that terminates in $O(1)$ iterations. For all iterations $t\in[\tau,T]$, $f(\theta_{t})\leq\frac{\lambda_{0}}{\lambda_{1}}$ and hence $\lambda_{0}\geq\lambda_{1}\,f(\theta_{t})$. Using this relation and following the standard proof for uniformly smooth functions satisfying the PL inequality completes the proof for Phase 2. which terminates in $O(\ln(1/\epsilon))$ iterations. Putting together the results for both phases completes the proof. ∎
The above result shows that GD-LS converges linearly, where the convergence rate depends on the ratio $\nicefrac{{\lambda_{0}}}{{\lambda_{1}}}$. On the other hand, for an $L$ uniformly-smooth function satisfying the PL condition, GD(1/L) requires $O\left(\frac{L}{\mu}\,\ln\left(\frac{f(\theta_{0})}{\epsilon}\right)\right)$ iterations (Karimi et al., 2016). Ignoring the constant first term independent of $\epsilon$ and assuming $\lambda_{0}\approx L$, we can see that the result in Theorem 3 is better than the standard rate when $\nicefrac{{\lambda_{0}}}{{\lambda_{1}}}$ is smaller than $f(\theta_{0})$. It is important to note that since GD-LS can automatically (without any change in the algorithm) exploit the uniform smoothness as well and obtain the standard result, the number of iterations required for GD-LS is $\min\left\{O\left(\frac{L}{\mu}\,\ln\left(\frac{f(\theta_{0})}{\epsilon}\right)\right),O\left(\frac{\lambda_{0}}{\mu}\,\ln\left(\frac{\lambda_{0}}{\lambda_{1}\,\epsilon}\right)\right)\right\}$, meaning that GD-LS should always converge at least as fast as GD(1/L). Since the GLM objective is also uniformly smooth (Mei et al., 2021, Lemma 10), we empirically verify our hypothesis in Fig. 3.
Figure 3: Comparing GD-LS with $c=1/2$, $\eta_{\max}=10^{4}$ and GD(1/L) for GLMs on a synthetic dataset with $n=10^{4}$, $d=200$, $\left\|\theta^{*}\right\|=1$. GD-LS converges faster than GD(1/L), and demonstrates Phase 1 behaviour initially, followed by a linear rate in Phase 2.
Comparing Theorem 3 for GLMs to the existing results, we note that Hazan et al. (2015, Lemma 3.1) show that the GLM objective is locally quasi-convex, and use this property to derive a slower $O(1/\epsilon^{2})$ convergence rate for normalized GD with a decreasing step-size (Hazan et al., 2015, Theorem 4.1). Finally, for the GLM objective, Mei et al. (2021) propose a novel variant of normalized GD and prove that it converges linearly, with a better constant dependence than GD. However, their method requires the knowledge of $\mu$, making it difficult to implement. On the other hand, GD-LS does not require parameter tuning and achieves similar theoretical results as these specialized methods.
@@ 6 SGD with Stochastic Line-search
In this section, we analyze the convergence of stochastic gradient descent (SGD) (Robbins & Monro, 1951) with a stochastic variant of the Armijo line-search (referred to as SGD-SLS) proposed in Vaswani et al. (2019b). We focus on the convex, finite-sum setting, and consider minimizing $f(\theta)=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\theta)$ where each $f_{i}$ is convex and satisfies Assumptions 1, 2 and 3. For ease of exposition, we assume that each $f_{i}$ is $(L_{0},L_{1})$ non-uniform smooth, and note that it is straightforward to analyze the case where the $f_{i}$’s have different smoothness constants. Binary classification using logistic regression and multi-class classification using the cross-entropy loss are examples of such an objective.
At iteration $t\in[T]$, SGD randomly samples a function ${f_{t}}$ from the $n$ functions in the finite-sum, computes its gradient and takes a descent step. Specifically,
$\theta_{t+1}=\theta_{t}-{\eta_{t}}\nabla{f_{t}}(\theta_{t})$ (8)
@@
Note that the stochastic Armijo condition only involves the sampled function and its gradient.
In order to analyze the convergence of SGD-SLS, we define $f_{i}^{*}:=\min f_{i}(\theta)$ as the minimum of function $i$ in the finite-sum and $\chi^{2}(\theta^{*}):=\mathbb{E}_{i}[f_{i}(\theta^{*})-f_{i}^{*}]$ as the “noise” in the stochastic gradients at the optimum (Loizou et al., 2021). In particular, if $\chi^{2}=0$, then each $f_{i}$ is minimized at $\theta^{*}$ implying that $\nabla f_{i}(\theta^{*})=0$. This special case is referred to as the interpolation setting (Vaswani et al., 2019a; Ma et al., 2018; Schmidt & Roux, 2013) and is useful in practical machine learning; for example, it is approximately satisfied by over-parameterized neural networks (Zhang et al., 2017) or non-parametric regression (Liang & Rakhlin, 2018; Belkin et al., 2019). Furthermore, logistic regression on linearly separable data is an example of a smooth convex loss that satisfies the interpolation condition and is the main motivation for the subsequent analysis.
When minimizing uniformly-smooth convex functions in the interpolation setting, Vaswani et al. (2019a) proved that SGD-SLS converges to the optimum at the $O(1/\epsilon)$ rate matching GD and faster than the standard $O(1/\epsilon^{2})$ rate (Bottou et al., 2018) for SGD. Moreover, SGD-SLS does not require the knowledge of the global smoothness constant, making it an attractive choice in practice. Motivated by our linear convergence results in Section 4, we analyze the convergence of SGD-SLS for convex functions that satisfy Assumptions 1, 2 and 3.
Since SLS involves an Armijo line-search for one (randomly chosen) function in each iteration, we can follow the same argument as in Lemma 1 and show that the step-size in each iteration is lower-bounded as ${\eta_{t}}\geq\min\left\{\eta_{\max},\frac{1}{\lambda_{0}\,+\lambda_{1}\,{f_{t}}(\theta_{t})}\right\}$. Given this result, we first prove the following lemma in Appendix E.
@@
For a fixed $\epsilon>0$, assuming $f(\theta):=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\theta)$ where each $f_{i}$ is convex, $L$ uniform smooth, satisfies Assumptions 1, 2 and 3 with $L_{0}=0,\,\omega=0$, if $\Delta_{t}:=\mathbb{E}[\left\|\theta_{t}-u\right\|_{2}^{2}]$, then, SGD-SLS guarantees that:
@@
$+2\eta_{\max}\,\chi^{2}(u)\,T$
where $u$ is an arbitrary comparator s.t. $f(u)<\mathbb{E}[f(\theta_{T})]$, $C:=\frac{(2c-1)}{c\,\lambda_{1}}$ and $\chi^{2}(u):=\mathbb{E}_{i}[f_{i}(u)-\min f_{i}(\theta)]$ is the noise in the stochastic gradients at $u$.
Since logistic regression on linearly separable data satisfies Assumptions 1, 2 and 3 and the interpolation condition, we follow a similar strategy as in Corollary 7 and prove the following corollary for logistic regression on separable data.
@@
For logistic regression on linearly separable data with margin $\gamma$, if, for all $i$, $\left\|x_{i}\right\|=1$, for a fixed $\epsilon\in\left(0,\frac{1}{8}\right)$, SGD-SLS with $\eta_{\max}=\frac{C}{\epsilon}$ where $C=\frac{(1-c)\,(2c-1)}{c\,\lambda_{1}}$ requires $T$ iterations to ensure that $\mathbb{E}[f(\theta_{T})]\leq 2\epsilon$ where,
$T\geq\frac{6\,c}{(1-c)\,(2c-1)\,\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}.$
Hence, SGD-SLS requires $O\left(\ln\left(\frac{1}{\epsilon}\right)^{2}\right)$ iterations (and hence gradient evaluations) to ensure convergence to an $\epsilon$ neighbourhood. This is an exponential improvement over the standard $O(1/\epsilon^{2})$ convergence rate for SGD on convex, uniformly smooth functions, where the improvements stem from effectively exploiting both interpolation and non-uniform smoothness.

### v2

@@ 5 GD-LS for Non-convex Losses
In this section, we consider non-convex losses that satisfy two alternative gradient domination conditions which enable convergence to the global optimum. In Sec. 5.1, we analyze the convergence of GD-LS for objectives corresponding to the softmax policy gradient in reinforcement learning. In Sec. 5.2, we consider objectives satisfying the Polyak-Łojasiewicz (PL) condition (Karimi et al., 2016; Polyak, 1987), and study the generalized linear model with a logistic link function as an example.
@@
$f$ satisfies a non-uniform gradient domination condition if there exists a constant $\zeta\geq 1$ and $\mu(\theta)>0$ s.t. for all $\theta$,
@@
Gradient domination or Łojasiewicz conditions are satisfied for matrix factorization (Ward & Kolda, 2023), policy gradient in reinforcement learning (Mei et al., 2020) and generalized linear models (Mei et al., 2021). These conditions have been exploited to prove global convergence guarantees for first-order methods (Karimi et al., 2016; Mei et al., 2021).
@@ 5.1 Gradient domination with $\zeta=1$
We use softmax policy optimization for multi-armed bandits (MAB) as a concrete example that satisfies Assn. 1, 2 and 3 and Assn. 4 with $\zeta=1$. In particular, we consider an MAB problem in the exact setting with deterministic, known rewards. This setting is often used as a testbed to evaluate policy gradient methods (Xiao, 2022; Mei et al., 2020; Lu et al., 2024). We prove the following proposition in App. A.
@@
where $a^{*}:=\operatorname*{arg\,max}_{a\in[K]}r(a)$ is the optimal arm. $f(\theta)$ is non-negative, satisfies Assn. 2 with $L_{0}=0$ and $L_{1}=72$Assn. 3 with $\nu=24$ and $\omega=0$ and Assn. 4 with $\zeta=1$ and $\mu(\theta)=\pi_{\theta}(a^{*})$.
Softmax policy gradient methods (Williams, 1992) optimize the above non-convex objective using gradient descent, and have been analyzed recently (Mei et al., 2020; Agarwal et al., 2021). We aim to use GD-LS to optimize the objective defined in Prop. 4 for which the GD update is given by: $\theta_{t+1}=\theta_{t}-{\eta_{t}}\,\nabla f(\theta_{t})=\theta_{t}+{\eta_{t}}\nabla_{\theta}\langle\pi_{\theta},r\rangle$, and the corresponding Armijo condition is equal to:
$\langle\pi_{\theta_{t+1}},r\rangle$ $\geq\langle\pi_{\theta},r\rangle+c{\eta_{t}}\,\left\|\nabla_{\theta}\langle\pi_{\theta},r\rangle\right\|_{2}^{2}.$
In Prop. 7 in App. A, we show that, under additional assumptions, the softmax policy gradient objective for tabular Markov decision processes (MDPs) also satisfies the Assn. 1, 2, 3 and 4 with $L_{0}=0$ $\omega=0$ and $\zeta=1$. In this case, the exact setting corresponds to knowing the rewards and the transition probabilities, and is the same setting under which classic RL algorithms such as value iteration or policy iteration are analyzed (Puterman, 2014).
We now characterize the convergence rate of GD-LS.
@@
For an $\epsilon>0$, assuming $f(\theta)$ satisfies Assn. 1, 2 and 3 with $L_{0}=0,\,\omega=0$ and Assn. 4 with $\zeta=1$, if $\mu:=\min_{t\in[T]}\mu(\theta_{t})$, then, GD-LS with $\eta_{\max}=\infty$, requires
@@
iterations to ensure $f(\theta_{T})\leq\epsilon$.
In order to better understand the implications of Cor. 3, we instantiate the above result for MAB and prove the following corollary in Sec. D.1.
@@
For an MAB problem with $K$ arms, rewards bounded in $[0,1]$, GD-LS with a uniform initialization i.e. afor-$\forall a$, $\pi_{\theta_{0}}(a)=\nicefrac{{1}}{{K}}$, $c=\frac{1}{2}$, $\eta_{\max}=\infty$ requires
$T=O\left(K^{2}\ln\left(\nicefrac{{1}}{{\epsilon}}\right)\right)$
iterations to guarantee $\langle\pi_{\theta_{T}},r\rangle\geq r(a^{*})-\epsilon$.
Hence, for MAB, GD-LS converges at a linear rate. Under additional assumptions, by combining Prop. 7 and 3, we can prove a similar linear rate for tabular MDPs in the exact setting (see Cor. 9 in Sec. D.1 for the formal result).
The above result is in contrast to GD(1/L) which can only attain an $\Omega\left(\frac{1}{\epsilon}\right)$ convergence rate for both bandits and MDPs (Mei et al., 2020, Theorem 9, 10). The convergence rate of GD-LS matches that of algorithms designed for this specific problem, including GD with a specific line-search that requires the knowledge of $f^{*}$ (Lu et al., 2024), GD with specific increasing step-sizes (Liu et al., 2024), normalized GD (Mei et al., 2021), natural policy gradient (Kakade & Langford, 2002; Xiao, 2022) and mirror descent with a log-sum-exp mirror map (Asad et al., 2024). From a practical perspective, Lu et al. (2024, Figure 1) empirically demonstrate the linear convergence of GD-LS on tabular Markov decision processes, and hence, Cor. 9 substantiates their results theoretically.
Two-layer neural networks: We note that Assn. 1, 2 and 3 and Assn. 4 with $\zeta=1$ are also satisfied when using the (i) exponential loss to train (ii) two-layer neural networks with a smoothed leaky-ReLU non-linearity and (iii) assuming that the training data is linearly separable (Taheri & Thrampoulidis, 2023, Lemmas 3,5). In this setting, Theorem 1 in Taheri & Thrampoulidis (2023) shows that normalized GD can result in linear convergence for the resulting non-convex objective. Since this problem also satisfies the required assumptions for Cor. 3, GD-LS also converges linearly and unlike normalized GD, it does not require knowing problem-dependent constants.
Hence, these results demonstrate the universality of GD-LS.
@@ 5.2 Gradient domination with $\zeta=2$
We use the generalized linear model (GLM) with a logistic link function as an example of an objective that satisfies the PL condition (which corresponds to Assn. 4 with $\zeta=2$).
@@
If $\sigma(\cdot)$ is the sigmoid function and $\pi_{i}(\theta):=\sigma(\langle x_{i},\theta\rangle)$, assuming that for all $i\in[n]$, $\left\|x\right\|_{i}\leq 1$, $y_{i}=\pi_{i}(\theta^{*})$ such that $\left\|\theta^{*}\right\|\leq D<\infty$ and $\upsilon(\theta):=\min_{i\in[n]}{\left\{\pi_{i}(\theta)\cdot\left(1-\pi_{i}(\theta)\right)\right\}}$, then the GLM objective in Eq. 4 satisfies Assn. 4 with $\zeta=2$ and $\mu(\theta)=64\,[\upsilon(\theta)]^{2}\,[\min\{\upsilon(\theta),\upsilon(\theta^{*})\}]^{2}$.
Similar to logistic regression, the PL constant $\mu$ depends on $\upsilon(\theta)$. However, unlike logistic regression where $y_{i}\in\{0,1\}$ and $\left\|\theta^{*}\right\|$ can be infinite for separable data, for GLMs, $y_{i}\in(0,1)$, $\left\|\theta^{*}\right\|$ is bounded and consequently, $\mu(\theta^{*})>0$. Hence, as long as $\left\|\theta_{t}\right\|<\infty$ for all iterates $t\in[T]$, $\mu(\theta)$ is bounded away from zero. However, we note that this does not preclude the case where the iterates initially diverge away from the solution, resulting in large $\left\|\theta_{t}\right\|$ and small (but non-zero) $\mu(\theta)$.
Recall that in Prop. 2, we have seen that the GLM objective satisfies Assn. 1, 2 and 3 with non-zero $L_{0},L_{1}$, $\nu,\omega$. Furthermore, since the targets $y_{i}=\pi_{i}(\theta^{*})$ are assumed to be realizable in Lemma 2, $f^{*}=0$. Given these considerations, we prove the following theorem in App. D and characterize the convergence of GD-LS for such an objective.
@@
For a fixed $\epsilon\in\left(0,\frac{\lambda_{0}}{\lambda_{1}}\right)$, if $f$ satisfies Assn. 1, 2 and 3 and Assn. 4 with $\zeta=2$, $f^{*}=0$ and if $\mu:=\min_{t\in[T]}\mu(\theta_{t})>0$, then, GD-LS with $\eta_{\max}=\infty$, requires
@@
The above result shows that GD-LS converges linearly, where the convergence rate depends on the ratio $\nicefrac{{\lambda_{0}}}{{\lambda_{1}}}$. On the other hand, for an $L$ uniformly-smooth function satisfying the PL condition, GD(1/L) requires $O\left(\frac{L}{\mu}\,\ln\left(\frac{f(\theta_{0})}{\epsilon}\right)\right)$ iterations (Karimi et al., 2016). Ignoring the constant first term which is independent of $\epsilon$ and assuming $\lambda_{0}\approx L$, we can see that the result in Thm. 3 is better than the standard rate when $\nicefrac{{\lambda_{0}}}{{\lambda_{1}}}$ is smaller than $f(\theta_{0})$. It is important to note that since GD-LS can automatically (without any change in the algorithm) exploit the uniform smoothness and obtain the standard result as well, the number of iterations required for GD-LS is $\min\left\{O\left(\frac{L}{\mu}\,\ln\left(\frac{f(\theta_{0})}{\epsilon}\right)\right),O\left(\frac{\lambda_{0}}{\mu}\,\ln\left(\frac{\lambda_{0}}{\lambda_{1}\,\epsilon}\right)\right)\right\}$, meaning that GD-LS converges at least as fast as GD(1/L). Since the GLM objective is also uniformly smooth (Mei et al., 2021, Lemma 10), in Fig. 3, we empirically compare to GD(1/L), and verify the faster convergence of GD-LS.
Figure 3: Comparing GD-LS with $c=1/2$, $\eta_{\max}=10^{4}$ and GD(1/L) for GLM on a synthetic dataset with $n=10^{4}$, $d=200$, $\left\|\theta^{*}\right\|=1$.
Comparing Thm. 3 to the existing results for the GLM objective, we note that Hazan et al. (2015, Lemma 3.1) show that the objective is locally quasi-convex, and use this property to derive a slower $O(1/\epsilon^{2})$ convergence rate for normalized GD with a decreasing step-size (Hazan et al., 2015, Theorem 4.1). On the other hand, Mei et al. (2021) propose a novel variant of normalized GD and prove that it converges linearly, with a better constant dependence compared to GD. However, their method requires the knowledge of $\mu$, making it difficult to implement. On the other hand, GD-LS does not require the knowledge of any problem-dependent constants, and achieves similar or better theoretical results compared to these specialized methods.
@@ 6 SGD with Stochastic Line-search
In this section, we analyze the convergence of stochastic gradient descent (SGD) (Robbins & Monro, 1951) with a stochastic variant of the Armijo line-search (referred to as SGD-SLS) proposed in Vaswani et al. (2019b).
We focus on the convex, finite-sum setting, and consider minimizing $f(\theta)=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\theta)$ where each $f_{i}$ is convex and satisfies Assn. 1, 2 and 3. Logistic regression and multi-class classification using the cross-entropy loss are examples of such an objective. For ease of exposition, we assume that each $f_{i}$ is $(L_{0},L_{1})$ non-uniform smooth, and note that it is straightforward to analyze the case where the $f_{i}$’s have different smoothness constants.
At iteration $t\in[T]$, SGD randomly samples a function ${f_{t}}$ from the $n$ functions in the finite-sum, computes its gradient and updates the parameters. Specifically,
$\theta_{t+1}=\theta_{t}-{\eta_{t}}\nabla{f_{t}}(\theta_{t})\,,$ (8)
@@
Note that the above stochastic Armijo condition only involves the sampled function and its gradient.
In order to analyze the convergence of SGD-SLS, we define $f_{i}^{*}:=\min f_{i}(\theta)$ as the minimum of function $i$ in the finite-sum and $\chi^{2}(\theta^{*}):=\mathbb{E}_{i}[f_{i}(\theta^{*})-f_{i}^{*}]$ as the measure of the stochasticity at the optimum (Loizou et al., 2021). In particular, if $\chi^{2}=0$, then each $f_{i}$ is minimized at $\theta^{*}$ implying that $\nabla f_{i}(\theta^{*})=0$. This special case is referred to as the interpolation setting (Vaswani et al., 2019a; Ma et al., 2018; Schmidt & Roux, 2013) and is useful in practical machine learning; for example, it is approximately satisfied by over-parameterized neural networks (Zhang et al., 2017) or non-parametric regression (Liang & Rakhlin, 2018; Belkin et al., 2019). Furthermore, logistic regression on linearly separable data is an example of a smooth convex loss that satisfies the interpolation condition and is the main motivation for the subsequent analysis.
When minimizing uniformly-smooth convex functions in the interpolation setting, Vaswani et al. (2019a) proved that SGD-SLS converges to the optimum at an $O(1/\epsilon)$ rate, matching GD and is faster than the standard $O(1/\epsilon^{2})$ rate for SGD (Bottou et al., 2018). Moreover, SGD-SLS does not require the knowledge of the global smoothness constant, making it an attractive choice in practice. Motivated by this and the results in Sec. 4, we analyze SGD-SLS for convex functions that satisfy Assn. 1, 2 and 3.
Since SLS involves an Armijo line-search for one (randomly chosen) function in each iteration, we can follow the same argument as in Lemma 1 and show that the step-size in each iteration is lower-bounded i.e. ${\eta_{t}}\geq\min\left\{\eta_{\max},\frac{1}{\lambda_{0}\,+\lambda_{1}\,{f_{t}}(\theta_{t})}\right\}$. Given this result, we first prove the following lemma in App. E.
@@
For a fixed $\epsilon>0$ and comparator $u$, assuming $f(\theta):=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\theta)$ where each $f_{i}$ is convex, satisfies Assn. 1, 2 and 3 with $L_{0}=0,\,\omega=0$, if $\Delta_{t}:=\mathbb{E}[\left\|\theta_{t}-u\right\|_{2}^{2}]$, $f(u)\leq f(\theta_{t})$ for all $t\in[T]$, and $C:=\frac{(2c-1)}{c\,\lambda_{1}}$, then, SGD-SLS guarantees that:
@@
$+2\eta_{\max}\,\chi^{2}(u)\,T\,,$
where the expectation is over the randomness in all iterations $t\in[T]$ and $\chi^{2}(u):=\mathbb{E}_{i}[f_{i}(u)-\min f_{i}(\theta)]$ is a measure of the stochasticity at $u$.
Since logistic regression on linearly separable data is convex, satisfies Assn. 1, 2 and 3 and the interpolation condition, we follow a similar strategy as in Cor. 2 and prove the following corollary.
@@
For logistic regression on linearly separable data with margin $\gamma$, if, for all $i$, $\left\|x_{i}\right\|\leq 1$, for a fixed $\epsilon\in\left(0,\frac{1}{8}\right)$, SGD-SLS with $\eta_{\max}=O\left(\frac{(1-c)\,(2c-1)}{c}\,\frac{1}{\epsilon}\right)$ requires
$T\geq O\left(\frac{c}{(1-c)\,(2c-1)\,\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}\right)$
iterations to ensure that $\mathbb{E}[\inf_{t\in[T]}f(\theta_{t})]\leq 4\epsilon$.
Hence, SGD-SLS requires $O\left(\ln\left(\frac{1}{\epsilon}\right)^{2}\right)$ iterations (and hence gradient evaluations) to ensure convergence to an $\epsilon$ neighbourhood. For logistic regression on separable data, this is an exponential improvement over the $O(1/\epsilon)$ rate for SGD-SLS in Vaswani et al. (2019b) and SGD with a Polyak step-size in Loizou et al. (2021). Compared to the deterministic result in Sec. 4 which holds for the last-iterate, the above result only holds for the best-iterate. Furthermore, while Cor. 2 required $\eta_{\max}=\infty$, the above result requires $\eta_{\max}=O(\nicefrac{{1}}{{\epsilon}})$ in order to handle the stochasticity.
From a practical perspective, Vaswani et al. (2019b); Galli et al. (2024) have empirically demonstrated the effectiveness of SGD-SLS and its non-monotone variant for convex losses corresponding to linear and logistic regression, and non-convex losses with deep neural networks. In particular, for logistic regression on the mushrooms dataset where interpolation is satisfied (Vaswani et al., 2019b, Figure 3)Cor. 5 explains the linear convergence of SGD-SLS.